OpenStudy (anonymous):

What is 5 pi over 3 radians in degrees? 60° 120° 210° 300°

OpenStudy (anonymous):

1pi =180 degrees 5pi/3 = 300 degrees

OpenStudy (anonymous):

thank you! do you mind helping me with another problem?

OpenStudy (anonymous):

Okay I can help you. :3

OpenStudy (anonymous):

While crossing the Atlantic, sailors spot two mermaids 120° apart on each end of an island that is 6 miles away. How far apart are the mermaids around the outer edge of the island to the nearest tenth of a mile? 12.6 miles 3.1 miles 7.2 miles 20.2 miles

OpenStudy (anonymous):

This is a triangle problems. so let me draw it for you

OpenStudy (anonymous):



OpenStudy (anonymous):

Does this help?

OpenStudy (anonymous):

yeah, that helped but I still don't know what the answer would be :( @BVB4rmy

OpenStudy (jdoe0001):

\(\bf degrees = \cfrac{{\color{red}{ radians}}\cdot 180}{\pi}\implies degrees = \cfrac{{\color{red}{ \frac{5\pi}{3}}}\cdot 180}{\pi}\)

OpenStudy (jdoe0001):

\(\bf degrees = \cfrac{{\color{red}{ radians}}\cdot 180}{\pi}\implies degrees = \cfrac{{\color{red}{ \frac{5\pi}{3}}}\cdot 180}{\pi} \\ \quad \\ \implies \cfrac{5\cancel{\pi} \cdot 60}{\cancel{\pi}}\)

OpenStudy (anonymous):

I'm still lost

OpenStudy (anonymous):

jdoe is trying to give you the equation to your first question.

OpenStudy (anonymous):

I'm done with the first question, I'm on the mermaid one now

OpenStudy (anonymous):

Have you learned Cos, Sin, and Tan

OpenStudy (jdoe0001):

ohh ahemm

OpenStudy (anonymous):

yes, but I have no idea how to do it and just need the answer :(

OpenStudy (anonymous):

Uhm well I know that they have given you the angle the opposite and the adjacent

OpenStudy (anonymous):

so I think it would be Tan(120) = x/6

OpenStudy (anonymous):

@hoblos

OpenStudy (jdoe0001):

ehhe

OpenStudy (anonymous):

@hoblos can you help?

OpenStudy (jdoe0001):

lemme see ahemm